how do I create a 3% sodium bicarbonate solution without using baking soda but by using a liquid that says sodium bicarbonate on it? so how do I make 3 percent out of it?
1
Expert's answer
2018-07-05T05:22:47-0400
nswer:
A three-percent solution of sodium bicarbonate can be prepared by dilution, from an existing solution for this you need to know the concentration. 3 % it is 30 gram of substance per 100 gram or ml of water. For example, initially we have 15/100. Need 3/100. Increase the ratio, so that out of 3 is 15: 3/100 = 15/500. The denominator received how much the whole new solution will be.
And from 500 we subtract 15 we get water as a result = 485. And in the beginning of water it was 100-10 = 85. It is clear that you need to add water 485 - 85 = 400. But for this it is necessary to know the initial concentration of the solution.